# Efficient Learning for Entropy-regularized Markov Decision Processes via Multilevel Monte Carlo

### v1

@@ 2.3 MLMC Estimators and Their Sample Complexities
This section utilizes the MLMC technique to design more sample efficient estimators. Specifically, observe that for any $n\in{\mathbb{N}}$, the iterate $Q_{n}$ defined by (2.1) admits the following telescoping decomposition:
@@ 2.3.1 Formulation of General MLMC Estimators
Definition 2.2 provides a unified framework for various stochastic approximations of the soft-Bellman operator considered in this paper. These approximations may be biased estimators due to the logarithm function in the soft-Bellman operator (2.2), and this approximation bias is reflected in the error bound of the MLMC estimator (Theorem 2.2). The variable $(K^{\theta})_{\theta\in\Theta}$ represents the number of samples $(A^{\theta})_{\theta\in\Theta}$ used to approximate the integral over $\mathcal{A}$, which can be either deterministic or stochastic. The biased plain MC estimator, as defined in Definition 2.1 in Section 2.3.3, corresponds to $K^{\theta}\equiv K\in{\mathbb{N}}^{*}$, while an unbiased estimator involving stochastic $(K^{\theta})_{\theta\in\Theta}$ will be introduced in Section 2.3.4.
@@
and define ${Q}^{\theta}_{n,M,\mathbf{T}}(s,a)$ by
$Q^{\theta}_{n,M,\mathbf{T}}(s,a)\coloneqq\rho_{\alpha}^{\beta}\left(\hat{Q}^{\theta}_{n,M,\mathbf{T}}(s,a)\right).$ (2.8)
@@
To implement the MLMC estimator recursively, let $Q_{0}$ be the initial guess for $Q^{\star}$, and $T_{\mathrm{approx}}$ be a procedure for approximating the soft-Bellman operator, utilizing samples drawn from the measure $\mu$ (Definition 2.2). Using $T_{\mathrm{approx}}$, we define the procedure $DT_{\mathrm{approx}}$ for approximating the difference of the soft-Bellman operator evaluated at two different Q-functions in (2.7), which applies $T_{\mathrm{approx}}$ to evaluate both Q-functions at the same state-action samples to ensure variance reduction (see Remark 2.4). The pseudocode for implementing the MLMC estimator is then presented in Algorithm 1.
@@ 2.3.2 Error Bounds of General MLMC Estimators
    There exists $L>0$, depending on $\alpha,\beta$ and $\tau$, such that for all bounded measurable functions $Q_{1},Q_{2}:\mathcal{S}\times\mathcal{A}\times\Omega\rightarrow{\mathbb{R}}$ and random variables $S:\Omega\to\mathcal{S}$ satisfying
@@
    $\|T^{\theta}Q_{1}(S)-T^{\theta}Q_{2}(S)\|_{L^{2}}\leq L\|Q_{1}(S,A^{(\theta,1)})-Q_{2}(S,A^{(\theta,1)})\|_{L^{2}},\quad\forall\theta\in\Theta.$ (2.9)
Assumption 3 states that the approximation operator $T^{\theta}$ preserves bounded functions, and is Lipschitz continuous in the $\|\cdot\|_{L^{2}}$ norm. These properties allow for controlling the rate of error propagation in the recursive construction of the MLMC estimator. One can easily show that the (exact) soft-Bellman operator satisfies Assumption 3 (see Lemmas 4.1 and 4.2). We show that both the biased plain Monte Carlo estimator (Section 2.3.3) and the unbiased estimator (Section 2.3.4) for the soft-Bellman operator satisfy Assumption 3.
@@
Again, notice that the error $E_{n,M,\mathbf{T}}$ is independent of $\theta$, since it only depends on the distributional properties of $Q^{\theta}_{n,M,\mathbf{T}}$. In particular, throughout the paper, we will often specialize the expressions only involving the distributional properties of $Q^{\theta}_{n,M,\mathbf{T}}$ (such as moments) by taking $\theta=0$.
@@
Suppose Assumptions 1, 2, and 3 hold. Let $(Q^{\theta}_{n,M,\mathbf{T}})_{\theta\in\Theta,n\in{\mathbb{N}},M\in{\mathbb{N}}^{*}}$ be the MLMC estimators given in Definition 2.3. Assume that $\gamma L<1$ with the constant $L$ in Assumption 3, and $M\in{\mathbb{N}}^{*}$ satisfies
@@
$\sigma_{\mathbf{T}}(s,a)\coloneqq\mathrm{Var}(T^{0}Q_{0}(S^{0}_{s,a}))^{\frac{1}{2}},\quad\delta_{\mathbf{T}}(s,a)\coloneqq\left|\mathbb{E}\left[T^{0}Q^{\star}(S^{0}_{s,a})-TQ^{\star}(S^{0}_{s,a})\right]\right|.$ (2.14)
@@ 2.3.3 Sample Complexity with a Plain Monte Carlo Approximation for $T$
$\Lambda_{M}\coloneqq\gamma L+\frac{1+2\gamma L}{\sqrt{M}}+\frac{\sqrt{\gamma}}{M^{1/4}}<1,\quad n\geq\frac{\log\varepsilon-\log D}{\log\Lambda_{M}},\quad K\geq 9\varepsilon^{-2}\left(\frac{\gamma L^{\prime}}{1-\Lambda_{M}}\right)^{2}$ (2.15)
with $D\coloneqq\frac{3}{2}\max\left(\beta-\alpha,2\gamma L^{\prime}\right)$ and $L^{\prime}\coloneqq\tau\left[L-1\right]$, the MLMCb estimator satisfies
$\sup_{s\in\mathcal{S},a\in\mathcal{A}}\|Q^{0}_{n,M,\mathbf{T}_{K}}(s,a)-Q^{\star}(s,a)\|_{L^{2}}\leq\varepsilon,$
@@
$K_{\varepsilon}$ $=\left\lceil\frac{9(\gamma L^{\prime})^{2}}{(1-\Lambda_{M_{0}})^{2}\varepsilon^{2}}\right\rceil,$
@@
$\mathfrak{C}_{n_{\varepsilon},M_{0},K_{\varepsilon}}\leq C\varepsilon^{-\frac{2}{\log\Lambda_{M_{0}}}\log\varepsilon-\kappa},$ (2.17)
@@
A comparison between Theorems 2.1 and 2.3 reveals that the MLMC technique achieves a quadratic reduction of the sample complexity of the simple iterative estimator, under the same model assumption. Indeed, observe that $\Lambda_{M_{0}}$ can be made arbitrarily close to $\gamma L$ by choosing a sufficiently large (but fixed) $M_{0}$. This along with (2.17) indicates that the MLMCb estimator can achieve an asymptotic complexity of order $\varepsilon^{\frac{-2\log\varepsilon}{\log\gamma L+\delta}(1+o(1))}$ as $\varepsilon\to 0$, for any sufficiently small $\delta>0$. Consequently, the MLMC method achieves a quadratic reduction in complexity, improving from the $\varepsilon^{\frac{-4\log\varepsilon}{\log\gamma L}}$ complexity of the simple iterative MC estimator in Theorem 2.1 to approximately $\varepsilon^{\frac{-2\log\varepsilon}{\log\gamma L}}$.
However, we note that the MLMC technique alone cannot achieve a polynomial complexity estimator due to the inherent bias of $\mathbf{T}_{K}$ in approximating the soft-Bellman operator $T$. As shown in Lemma 4.4, the bias of $\mathbf{T}_{K}$ is of magnitude $\mathcal{O}(K^{-1/2})$. According to the error bound (2.13), achieving an accuracy $\varepsilon>0$ requires the number of fixed-point iterations to be $n_{\varepsilon}=\mathcal{O}(\log(\varepsilon^{-1}))$, while the sample size $K_{\varepsilon}$ diverges to infinity as $\varepsilon\to 0$. As a result, the total complexity scales as $K_{\varepsilon}^{n_{\varepsilon}}=\varepsilon^{-\log K_{\varepsilon}}$, indicating a super-polynomial rate as $\varepsilon\to 0$.
@@ 2.3.4 Sample Complexity with an unbiased Approximation for $T$
where for all $n$, $\overline{X}_{n}\coloneqq\frac{1}{n}\sum_{i=1}^{n}X_{i}$ and $p_{n}>0$ is is any sequence satisfying $\sum_{i=1}^{\infty}p_{i}=1$. This indicates that given an independent random variable $N$ with $\mathbb{P}(N=n)=p_{n}$ for all $n$, the estimator $Y\coloneqq p_{N}^{-1}\left(g(\overline{X}_{N+1})-g(\overline{X}_{N})\right)+g(\overline{X}_{1})$ is an unbiased estimator of $g(\mathbb{E}X)$. [4] further refines this estimator by using antithetic variates and a random sample size $N=2^{K}$, where $K$ follows a suitably chosen geometric distribution; see [5, 32] for related ideas.
@@
$\Delta_{K}^{\theta}Q(s)=g\left(\frac{1}{2^{K+1}}\sum_{k=1}^{2^{K+1}}\exp\left(-Q(s,A^{(\theta,k)})/\tau\right)\right)$
$-\frac{1}{2}\left[g\left(\frac{1}{2^{K}}\sum_{k=1}^{2^{K}}\exp\left(-Q(s,A^{(\theta,2k)})/\tau\right)\right)+g\left(\frac{1}{2^{K}}\sum_{k=1}^{2^{K}}\exp\left(-Q(s,A^{(\theta,2k-1)})/\tau\right)\right)\right],$
@@
We denote by $({Q}_{n,M,\tilde{\mathbf{T}}}^{\theta})_{n\in{\mathbb{N}},M\in{\mathbb{N}}^{*},\theta\in\Theta}$ the MLMC estimators defined using $\tilde{\mathbf{T}}=(\tilde{T}^{\theta})_{\theta\in\Theta}$ (Definition 2.3). These estimators will be referred to as MLMC estimators with unbiased estimation, or, in abbreviation, the MLMCu estimators.
@@
The parameter $r$ for the geometric distribution determines both the sample complexity and the stability of the Blanchet–Glynn approximation $\tilde{T}^{\theta}$. Observe that the expected sample size of $\tilde{T}^{\theta}$ is $\mathbb{E}[2^{\tilde{K}^{\theta}+1}]=\sum_{n=0}^{\infty}2^{n+1}p(n)=2r\sum_{n=0}^{\infty}2^{n}(1-r)^{n}=(2r-1)^{-1}2r$, and hence $\tilde{T}^{\theta}$ has a finite expected sample size. The condition $r<3/4$ ensures that the approximation $\tilde{T}^{\theta}Q$ has a finite variance for any given $Q$ (Proposition 2.4), and the map $Q\mapsto\tilde{T}^{\theta}Q$ is Lipschitz continuous with respect to the $\|\cdot\|_{L^{2}}$-norm (Proposition 2.5).
@@
It is easy to see that the family $\tilde{\mathbf{T}}$ is admissible in the sense of Definition 2.2. The following proposition shows that $\tilde{T}^{\theta}$ is unbiased and has a finite variance, whose proof is given in Section 5.8.
@@
Suppose Assumptions 1 and 2 hold. Let $({Q}_{n,M,\tilde{\mathbf{T}}}^{\theta})_{n\in{\mathbb{N}},M\in{\mathbb{N}}^{*},\theta\in\Theta}$ be the MLMCu estimator defined in Definition 2.5, and $L=L_{(\alpha,\beta,\tau,r)}$ be the Lipschitz constant in Proposition 2.5. Assume that $\gamma L<1$. Then for all $\varepsilon>0$, by setting $n,M\in{\mathbb{N}}^{*}$ such that
@@
$\sup_{s\in\mathcal{S},a\in\mathcal{A}}\|Q^{0}_{n,M,\tilde{\mathbf{T}}}(s,a)-Q^{\star}(s,a)\|_{L^{2}}\leq\varepsilon,$ (2.22)
@@
The condition $\gamma L<1$ in Theorem 2.6 involves a different constant $L$ than in Theorems 2.1 and 2.3. This condition holds if the discount factor $\gamma$ is sufficiently small. Indeed, observe that the Lipschitz constant $L$ depends on $\gamma$ only through $\alpha=(1-\gamma)^{-1}c_{\min}$ and $\beta=(1-\gamma)^{-1}c_{\max}$. As $\alpha$ and $\beta$ remain bounded as $\gamma\to 0$, and the map $(\alpha,\beta)\mapsto L_{(\alpha,\beta,\tau,r)}$ is continuous, it follows that $L$ remains bounded as $\gamma\to 0$ and $\lim_{\gamma\to 0}\gamma L=0$. However, it is unclear whether the condition $\gamma L<1$ holds for sufficiently large regularization parameter $\tau$.

### v3

@@ 2.3 MLMC Estimators and Their Sample Complexities
This section uses the MLMC technique to design more sample efficient estimators. Specifically, observe that for any $n\in{\mathbb{N}}$, the iterate $Q_{n}$ defined by (2.1) admits the following telescoping decomposition:
@@ 2.3.1 Formulation of General MLMC Estimators
Definition 2.2 provides a unified framework for various stochastic approximations of the soft-Bellman operator considered in this paper. These approximations may be biased estimators due to the logarithm function in the soft-Bellman operator (2.2), and this approximation bias is reflected in the error bound of the MLMC estimator (Theorem 2.2). The variable $(K^{\theta})_{\theta\in\Theta}$ represents the number of samples $(A^{\theta})_{\theta\in\Theta}$ used to approximate the integral over $\mathcal{A}$, which can be either deterministic or stochastic. The biased plain MC estimator, as defined in Definition 2.1, corresponds to $K^{\theta}\equiv K\in{\mathbb{N}}^{*}$, while an unbiased estimator involving stochastic $(K^{\theta})_{\theta\in\Theta}$ will be introduced in Section 2.3.4.
@@
and define ${Q}^{\theta}_{n,M,\mathbf{T}}(s,a)$ by $Q^{\theta}_{n,M,\mathbf{T}}(s,a)\coloneqq\rho_{\alpha}^{\beta}\left(\hat{Q}^{\theta}_{n,M,\mathbf{T}}(s,a)\right).$
@@
To implement the MLMC estimator recursively, let $Q_{0}$ be the initial guess for $Q^{\star}$, and $T_{\mathrm{approx}}$ be a procedure for approximating the soft-Bellman operator, using samples drawn from the measure $\mu$ (cf. Definition 2.2). Using $T_{\mathrm{approx}}$, we define the procedure $DT_{\mathrm{approx}}$ for approximating the difference of the soft-Bellman operator evaluated at two different Q-functions in (2.7), which applies $T_{\mathrm{approx}}$ to evaluate both Q-functions at the same state-action samples to ensure variance reduction (see Remark 2.4). The pseudocode for implementing the MLMC estimator is then presented in Algorithm 1.
@@ 2.3.2 Error Bounds of General MLMC Estimators
    There exists $L>0$, depending on $\alpha,\beta$ and $\tau$, such that for all bounded measurable functions $Q_{1},Q_{2}:\mathcal{S}\times\mathcal{A}\times\Omega\rightarrow{\mathbb{R}}$ and random variable $S:\Omega\to\mathcal{S}$ satisfying
@@
    $\left\|T^{\theta}Q_{1}(S)-T^{\theta}Q_{2}(S)\right\|_{L^{2}}\leq L\left\|Q_{1}\left(S,A^{(\theta,1)}\right)-Q_{2}\left(S,A^{(\theta,1)}\right)\right\|_{L^{2}},\quad\forall\theta\in\Theta.$ (2.8)
Assumption 3 states that the approximation operator $T^{\theta}$ preserves bounded functions when taking expectation, and is Lipschitz continuous in the $\|\cdot\|_{L^{2}}$ norm. These properties allow for controlling the rate of error propagation in the recursive construction of the MLMC estimator. One can easily show that the (exact) soft-Bellman operator satisfies Assumption 3 (see Lemmas 4.1 and 4.2). We show that both the biased plain Monte Carlo estimator (Lemma 4.3) and the unbiased estimator (Proposition 2.5) for the soft-Bellman operator satisfy Assumption 3.
@@
Notice that the error $E_{n,M,\mathbf{T}}$ is independent of $\theta$, since it only depends on the distributional properties of $Q^{\theta}_{n,M,\mathbf{T}}$. In particular, throughout the paper, we will often specialize the expressions only involving the distributional properties of $Q^{\theta}_{n,M,\mathbf{T}}$ (such as moments) by taking $\theta=0$.
@@
Suppose Assumptions 1, 2, and 3 hold. Let $\left(Q^{\theta}_{n,M,\mathbf{T}}\right)_{\theta\in\Theta,n\in{\mathbb{N}},M\in{\mathbb{N}}^{*}}$ be the MLMC estimators given in Definition 2.3. Assume that $\gamma L<1$ with the constant $L$ in Assumption 3, and $M\in{\mathbb{N}}^{*}$ satisfies
@@
$\sigma_{\mathbf{T}}(s,a)\coloneqq\mathrm{Var}\left(T^{0}Q_{0}\left(S^{0}_{s,a}\right)\right)^{\frac{1}{2}},\quad\delta_{\mathbf{T}}(s,a)\coloneqq\left|\mathbb{E}\left[T^{0}Q^{\star}\left(S^{0}_{s,a}\right)-TQ^{\star}\left(S^{0}_{s,a}\right)\right]\right|.$ (2.13)
@@ 2.3.3 Sample Complexity with a Plain Monte Carlo Approximation for $T$
$\Lambda_{M}\coloneqq\gamma L+\frac{1+2\gamma L}{\sqrt{M}}+\frac{\sqrt{\gamma}}{M^{1/4}}<1,\quad n\geq\frac{\log\varepsilon-\log D}{\log\Lambda_{M}},\quad K\geq\frac{3\gamma(L^{\prime})^{2}}{2\tau(1-\Lambda_{M})\varepsilon}$ (2.14)
with $D\coloneqq\frac{3}{2}\max\left(\beta-\alpha,2\gamma L^{\prime}\right)$ and $L^{\prime}\coloneqq\tau(L-1)$, the MLMCb estimator satisfies
$\sup_{s\in\mathcal{S},a\in\mathcal{A}}\left\|Q^{0}_{n,M,\mathbf{T}_{K}}(s,a)-Q^{\star}(s,a)\right\|_{L^{2}}\leq\varepsilon,$
@@
$K_{\varepsilon}$ $=\left\lceil\frac{3\gamma(L^{\prime})^{2}}{2\tau(1-\Lambda_{M_{0}})\varepsilon}\right\rceil,$
@@
$\mathfrak{C}_{n_{\varepsilon},M_{0},K_{\varepsilon}}\leq C\varepsilon^{-\frac{\log\varepsilon}{\log\Lambda_{M_{0}}}-\kappa},$ (2.16)
@@
A comparison between Theorems 2.1 and 2.3 reveals that the MLMC technique achieves a cubic reduction of the sample complexity of the simple iterative estimator, under the same model assumption. Indeed, observe that $\Lambda_{M_{0}}$ can be made arbitrarily close to $\gamma L$ by choosing a sufficiently large (but fixed) $M_{0}$. This along with (2.16) indicates that the MLMCb estimator can achieve an asymptotic complexity of order $\varepsilon^{\frac{-\log\varepsilon}{\log\gamma L+\delta}(1+o(1))}$ as $\varepsilon\to 0$, for any sufficiently small $\delta>0$. Consequently, the MLMC method achieves a cubic reduction in complexity, improving from the $\varepsilon^{\frac{-3\log\varepsilon}{\log\gamma L}}$ complexity of the simple iterative MC estimator in Theorem 2.1 to approximately $\varepsilon^{\frac{-\log\varepsilon}{\log\gamma L}}$.
However, we note that the MLMC technique alone cannot achieve a polynomial complexity estimator due to the inherent bias of $\mathbf{T}_{K}$ in approximating the soft-Bellman operator $T$. As shown in Lemma 4.4, the bias of $\mathbf{T}_{K}$ is of magnitude $\mathcal{O}(K^{-1})$. According to the error bound (2.12), achieving an accuracy $\varepsilon>0$ requires the number of fixed-point iterations to be $n_{\varepsilon}=\mathcal{O}(\log(\varepsilon^{-1}))$, while the sample size $K_{\varepsilon}$ diverges to infinity as $\varepsilon\to 0$. As a result, the total complexity scales as $K_{\varepsilon}^{n_{\varepsilon}}=\varepsilon^{-\log K_{\varepsilon}}$, indicating a super-polynomial rate as $\varepsilon\to 0$.
@@ 2.3.4 Sample Complexity with an unbiased Approximation for $T$
where for all $n$, $\overline{X}_{n}\coloneqq\frac{1}{n}\sum_{i=1}^{n}X_{i}$ and $p_{n}>0$ is is any sequence satisfying $\sum_{i=1}^{\infty}p_{i}=1$. This indicates that given an independent random variable $N$ with $\mathbb{P}(N=n)=p_{n}$ for all $n$, the estimator $Y\coloneqq p_{N}^{-1}\left(g(\overline{X}_{N+1})-g(\overline{X}_{N})\right)+g(\overline{X}_{1})$ is an unbiased estimator of $g(\mathbb{E}X)$. [4] further refine this estimator by using antithetic variates and a random sample size $N=2^{K}$, where $K$ follows a suitably chosen geometric distribution; see the work of [5, 33] for related ideas.
@@
$\Delta_{K}^{\theta}Q(s)=g\left(\frac{1}{2^{K+1}}\sum_{k=1}^{2^{K+1}}\exp\left(-Q\left(s,A^{(\theta,k)}\right)/\tau\right)\right)$
$-\frac{1}{2}\left[g\left(\frac{1}{2^{K}}\sum_{k=1}^{2^{K}}\exp\left(-Q\left(s,A^{(\theta,2k)}\right)/\tau\right)\right)+g\left(\frac{1}{2^{K}}\sum_{k=1}^{2^{K}}\exp\left(-Q\left(s,A^{(\theta,2k-1)}\right)/\tau\right)\right)\right],$
@@
We denote by $({Q}_{n,M,\tilde{\mathbf{T}}}^{\theta})_{n\in{\mathbb{N}},M\in{\mathbb{N}}^{*},\theta\in\Theta}$ the MLMC estimators defined using $\tilde{\mathbf{T}}=(\tilde{T}^{\theta})_{\theta\in\Theta}$ (cf. Definition 2.3). These estimators will be referred to as MLMC estimators with unbiased estimation, or, in abbreviation, the MLMCu estimators.
@@
The parameter $r$ for the geometric distribution determines both the sample complexity and the stability of the Blanchet–Glynn approximation $\tilde{T}^{\theta}$. Observe that the expected sample size of $\tilde{T}^{\theta}$ is $\mathbb{E}[2^{\tilde{K}^{\theta}+1}]=\sum_{n=0}^{\infty}2^{n+1}p(n)=2r\sum_{n=0}^{\infty}2^{n}(1-r)^{n}=(2r-1)^{-1}2r$, which is finite for $r>1/2$. The condition $r<3/4$ ensures that the approximation $\tilde{T}^{\theta}Q$ has a finite variance for any given $Q$ (Proposition 2.4), and the map $Q\mapsto\tilde{T}^{\theta}Q$ is Lipschitz continuous with respect to the $\|\cdot\|_{L^{2}}$-norm (Proposition 2.5).
@@
It is easy to see that the family $\tilde{\mathbf{T}}$ is admissible in the sense of Definition 2.2. The following proposition shows that $\tilde{T}^{\theta}$ is unbiased and has finite variance. The proof is given in Section 5.8.
@@
Suppose Assumptions 1 and 2 hold. Let $\left({Q}_{n,M,\tilde{\mathbf{T}}}^{\theta}\right)_{n\in{\mathbb{N}},M\in{\mathbb{N}}^{*},\theta\in\Theta}$ be the estimators defined in Definition 2.5, and $L=L_{(\alpha,\beta,\tau,r)}$ be the Lipschitz constant in Proposition 2.5. Assume that $\gamma L<1$. Then for all $\varepsilon>0$, by setting $n,M\in{\mathbb{N}}^{*}$ such that
@@
$\sup_{s\in\mathcal{S},a\in\mathcal{A}}\left\|Q^{0}_{n,M,\tilde{\mathbf{T}}}(s,a)-Q^{\star}(s,a)\right\|_{L^{2}}\leq\varepsilon,$ (2.21)
@@
The condition $\gamma L<1$ in Theorem 2.6 involves a different constant $L$ than in Theorems 2.1 and 2.3. This condition holds if the discount factor $\gamma$ is sufficiently small. Indeed, observe that the Lipschitz constant $L$ depends on $\gamma$ only through $\alpha=(1-\gamma)^{-1}c_{\min}$ and $\beta=(1-\gamma)^{-1}c_{\max}$. As $\alpha$ and $\beta$ remain bounded as $\gamma\to 0$, and the map $(\alpha,\beta)\mapsto L_{(\alpha,\beta,\tau,r)}$ is continuous, it follows that $L$ is bounded as $\gamma\to 0$ and $\lim_{\gamma\to 0}\gamma L=0$. However, it is unclear whether $\gamma L<1$ holds for a sufficiently large regularization parameter $\tau$.